OpenStudy (anonymous):

The number N of employees at the Ajex Steel Company is growing at a rate given by the equation N'(t) = 20e^(.01t) where N is the number of people employed at time t, in years. The number of employees currently is 400. Find an equation for N(t).

hero (hero):

If no one else posts a solution I will. Give me a minute

OpenStudy (anonymous):

oh thank u

OpenStudy (anonymous):

umm i have the solution i just dont know how to set it up

hero (hero):

I understand

hero (hero):

1. To find N(t), take the integral of N'(t): \[N(t) = \int\limits{ 20e^{.01t} dt}\]2. Remove the constant, 20, from the integral: \[20 \int\limits\limits{ e^{.01t} dt}\]3. Let u = .01t Then du = .01dt and: \[dt = \frac{du}{.01}\]4. Next substitute everything in terms of u and du: \[20 \int\limits\limits\limits{ e^{u} \frac{du}{.01}}\]5. Move the constant 1/.01 outside of the integal: \[\frac{20}{.01} \int\limits\limits\limits\limits{ e^{u} du}\] Can you finish the rest from here?

OpenStudy (anonymous):

yes thank you

hero (hero):

Don't forget to add plus C after taking the integral!

OpenStudy (anonymous):

oh yes

OpenStudy (mimi_x3):

\[20*\frac{1}{0.01} \int\limits e^{0.01t} dt\]

OpenStudy (anonymous):

ok i got the part N(t) = 2000e^(.01t) + C but the solution says its N(t)=2000e^(.01t) - 1600? what am i doing wrong?

OpenStudy (mimi_x3):

did you find the constant yet?

hero (hero):

Yup that's what has to be found

OpenStudy (anonymous):

how do you do that?

OpenStudy (mimi_x3):

400 = 2000^(0.01t) + C i think..

OpenStudy (mimi_x3):

wait..what's the point of me being here since hero know's how to do it :/

OpenStudy (mimi_x3):

Hero, i think that there is something wrong..

hero (hero):

Nothing is wrong. If it was, the first part would have been incorrect. The constant is what needs to be found.

OpenStudy (mimi_x3):

okay, im out; since you know how to do it..

hero (hero):

Thanks for your contribution @Mimi_x3 . Of course it actually looks like you're bailing out on us but that's okay.

OpenStudy (anonymous):

lol

OpenStudy (mimi_x3):

it's 1am for godsakes.. to find the constant.. 400 = 2000e^(0.01t) + C

OpenStudy (mimi_x3):

400 = 2000e^(0.01*0) + C C = -16000 therefore; 2000e^(0.01t) -16000 Done!

OpenStudy (mimi_x3):

-1600*

OpenStudy (mimi_x3):

happy now?

hero (hero):

Go to bed now :P

OpenStudy (anonymous):

lol thanks

OpenStudy (mimi_x3):

wait..you did not know how to do this hero?

hero (hero):

I did actually.

OpenStudy (anonymous):

idk who to give the medal to

OpenStudy (mimi_x3):

oh gawd..waster my time arghh

OpenStudy (anonymous):

lol

OpenStudy (mimi_x3):

give it to hero..i dont need it

hero (hero):

@myluv4u , I already gave @Mimi_x3 a medal, but I am the one who helped you initially.

OpenStudy (anonymous):

well thank you both for the help

OpenStudy (anonymous):

thank you hero i had already gave you the medal , i forgot